OpenStudy (imtant):

Why e^In y=e^2x^2+c become y=e^2x^2*e^c?

OpenStudy (zzr0ck3r):

it looks like what you had was \[ln(y) = 2x^2+c\] and then to solve for y, we raise both sides as a power of e \[e^{ln(y)}=e^{2x^2+c}\] now on the right had side we use the property that \[a^{b+c}=a^b*a^c\] and on the left hand side we use the fact that \[e^{ln(y)}=y\]putting those both together we have \[y=e^{2x^2+c}=e^{2x^2}*e^c\] does this make sense?

OpenStudy (zzr0ck3r):

so\[e^{ln(y)}=e^{2x^2+c}\iff y=e^{2x^2}e^c\]

OpenStudy (imtant):

Okay, I got it. Thanks zzrock3r. :)